OpenStudy (anonymous):

Find r, given the following: A = P(1+r/n)^nt....and: 10550 = 5150 (1 + r/360)^360*19. #tkhunny

OpenStudy (anonymous):

In short, I just need to find out how to create a new equation from this one.

OpenStudy (anonymous):

I've tried once, but to no avail (i.e. I got the wrong answer)

OpenStudy (tkhunny):

Divide by 5150 2.0485436893203883495145631067961 = (1 + r/360)^{6840} Find the 6840th root (raise to the 1/6840 power). You could also introduce logarithms. 1.0001048489385532660135203782234 = 1 + r/360 Can you finish.

OpenStudy (anonymous):

Right, the first step kinda goes without saying. Wait... just tried this: ((1.0001048489385532660135203782234)^(1/6840)-1)*360

OpenStudy (anonymous):

Which is apparently wrong. So, in short, no, I'm sorry, but I can't finish the problem myself.

OpenStudy (tkhunny):

No, that's crazy. Where dod you even get that? Subtract 1 1.0484893855326601352037822344391e-4 = r/360 Multiply by 360 r = 0.03774561787917576486733616043981 Sorry, I was just dumping decimal places. I'll let you worry about thr rounding rules.

OpenStudy (anonymous):

Right...I also tried that, actually...but it didn't accept that answer either....*_*

OpenStudy (anonymous):

Okay...I only have one attempt left...I give up on this one. Thank you for your help though!